Solve the first order linear inhomogeneous differential equation using the constant variation method

xy,-2y=2x4

1
Expert's answer
2022-02-07T16:23:36-0500

Given a first order inhomogeneous linear differential equation of the form;

"\displaystyle
y\prime+p(x)y=f(x)\cdots\cdots\cdots\color{red}(1)"

using constant variation method, the general solution is given by;

"\displaystyle
y(x)=v(x)e^{q(x)}+ce^{q(x)}"

where "\displaystyle
v\prime(x)=e^{-q(x)}f(x)", "\displaystyle
q(x)=\int[-p(x)]\ dx", and "c" is an arbitrary constant.


Now, the given DE is of the form;

"y\prime-\frac{2}{x}y=2x^3"

By comparing the given DE with "\color{red}(1)";

"\displaystyle
q(x)=\int[-p(x)]\ dx=\int\frac{2}{x}\ dx=2\ln x=\ln x^2"

"\displaystyle
v\prime(x)=e^{-q(x)\ dx}\times f(x)=e^{-\ln x^2}\times2x^3=\frac{1}{x^2}\times2x^3=2x"

"\displaystyle
\Rightarrow v(x)=\int2x\ dx=x^2"


Thus, the general solution is;

"\displaystyle
y(x)=v(x)e^{q(x)}+ce^{q(x)}=x^2e^{\ln x^2}+ce^{\ln{x^2}}=x^4+cx^2"